OpenStudy (anonymous):

Math Help Please!!!! Write an equation of the direct variation that includes the point (9, -12). A) y = 1 1/3x B) y = -1 1/3x C)y = -3/4x D) y = -1/12x

OpenStudy (godlovesme):

y=kx -12= k9 solve for k @Answers101

OpenStudy (godlovesme):

@Answers101 u there?

OpenStudy (anonymous):

Yes

OpenStudy (godlovesme):

so what's -12/9 ?

OpenStudy (xapproachesinfinity):

remember what i wrote in the other post \[\large y=\frac{k}{x}\] if the point lies on the graph of the equation then we just substitute x and y by (9, -12) to find k so \[\Large -12=\frac{k}{9} \Longrightarrow k=-12\times9=-108\]

OpenStudy (anonymous):

1 1/3 @Godlovesme

OpenStudy (godlovesme):

i think u forgot the -ve sign @Answers101

OpenStudy (xapproachesinfinity):

oops we are looking for a direct variation i did the opposite \[y=kx \Longrightarrow -12=9k \Longrightarrow k=\frac{-12}{9}=\frac{-4}{3}\] so the final answer is \[y=\frac{-4}{3}x\]

OpenStudy (xapproachesinfinity):

pls make sure what you post errors in your choices ????

OpenStudy (xapproachesinfinity):

@Answers101 see here!!